OpenStudy (anonymous):

SAT Question A box contains only red pens and white pens. If one red pen is removed at random the probability that it is red is 1/5. After putting another 600 red pens in the box, the probability of selecting a red pen is 4/5. How many pens were originally in the box? Can someone explain the steps? (Answer is 200)

OpenStudy (anonymous):

iam uncertain about the method but i know why the ans is correct

OpenStudy (anonymous):

Let \(r\) and \(w\) be the number of red and white pens, respectively. In any given draw, the probability of getting a red pen would be \(\dfrac{r}{r+w}\). Say the number of pens in the first draw is \(P_1\), and the second draw has \(P_2\) pens. \[\frac{r}{P_1}=\frac{1}{5}\] \[\frac{r}{P_2}=\frac{4}{5}~\text{ where }P_2=P_1+600\] Basically, you have a system of two equations with two unknowns. You should be able to solve for \(r\) and \(P_1\), I think.

OpenStudy (anonymous):

Hmm, apparently that system has negative solutions... weird

OpenStudy (anonymous):

Ah, got it, it should be \[\begin{cases} \dfrac{r}{P_1}=\dfrac{1}{5}\\\\ \dfrac{r+600}{P_1+600}=\dfrac{4}{5} \end{cases}\]

OpenStudy (anonymous):

Do I just cross multiply your equations?

OpenStudy (anonymous):

Well I would first establish a relation between \(r\) and \(P_1\), then substitute into the second equation.

OpenStudy (anonymous):

Cross-multiplication will happen when you solve for the substituted variable.

OpenStudy (aum):

Let 'r' be the number of red pins and 'w' be the number of white pens at the start. \(\Large \frac{r}{r+w} = \frac 15;~~~~ \frac{r+600}{r+600+w} = \frac 45 \) Eliminate 'w' and solve for 'r' From (1) : 5r = r + w w = 4r From (2): 5r + 3000 = 4r + 2400 + 4w r + 600 = 4w put w = 4r r + 600 = 16r 15r = 600 r = 40 w = 4r = 160 r + w = 40 + 160 = 200 (Assumption: The first pen that was drawn was put back).

OpenStudy (anonymous):

^what @aum said :)

OpenStudy (anonymous):

Ah I got confused around r + 600 = 4w put w = 4r This part. Thanks for your help guys.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Wish I could give two medals.

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

letz do 1 one thing u give to @aum i will giv to @SithsAndGiggles

OpenStudy (anonymous):

thanks both of u for sharing ur knowledge

OpenStudy (aum):

you are welcome.

OpenStudy (anonymous):

Everyone's welcome